# Distinguishing Ignorance from Error in LLM Hallucinations

### v1

@@ Appendix A Hyper parameters search for Knowledge Categorization
Knowledge detection typically relies on the model’s output, either through logits or generation. We focus on the generation approach, assessing whether the model consistently produces a factually correct answer among multiple samples, similar to recent work [Gekhman et al., 2024]. This method is influenced by various hyperparameters including (1) number of generations, (2) sampling temperature, (3) length of generation, and (4) prompt structure.
@@
We tested this approach on 1000 random TriviaQA examples across our three models. The average similarity among all 8 configurations (28 unique combinations) was $93.6\%$ for Llama, $92.7\%$ for Mistral, and $92.2\%$ for Gemma, indicating a high consistency in knowledge classifications. The lowest similarity (about $80\%$) occurred with zero-shot configurations.
Based on these results, we adopted the baseline setting as our knowledge detector, using the 3-shot prompt corresponding to the bad-shots used in subsequent hallucination classification. The high similarity between different few-shot prompts suggests that varying the few-shot examples should yield comparable results. To enhance reliability, we supplemented this approach with one greedy generation, ensuring we capture the most likely output even if temperature-based generations fail to produce it.
@@
Appendix B Dataset construction specifics
@@
For the Natural Questions datasaet, we also used 30K random examples, excluding examples with answer longer than 5 tokens as well as examples without an answer or with more than one answer. We again added lower-case versions of upper-case answers.
@@ B.1 Generic dataset construction
To create the generic dataset, the key addition was obtaining an incorrect answer for each example. We generated these using Mistral with the following prompt:
@@
Table 3: Examples from the generic dataset of TriviaQA.
@@ B.2 Specific dataset construction
Subsequently, we employed a prompt-based approach utilizing bad-shots to generate responses from the model. For each example, the model produced a greedy generation of 5 tokens. We then evaluated these generations to categorize the examples: if the correct answer was present within the generated tokens, we labeled the example as “factually correct”; otherwise, it was labeled as “hallucinated”.
@@
Appendix C Alice-Bob-setting dataset statistics
We show in Table 4 the dataset statistics for the Alice-Bob setting. There are fewer hallucinations than in the bad-shot setting; however, there are still at least 600 hallucinations in each of the configurations, enabling sufficient investigation.
Table 4: Dataset label statistics on the Alice-Bob setting.
@@ Appendix C Alice-Bob-setting dataset statistics
Appendix D Detection results on the MLP and Attention components
The results are shown in Figures 8 and 9 for the classification into the two hallucination types and factually correct examples, for the bad-shot and Alice-Bob settings.
Next in Figures 10 and 11 we see the results of the generalization of the bad-shot setting to the Alice-Bob setting using the MLP and Attention components.
Lastly, Figures 12 and 13 show detection of HK+ hallucination results using classifiers trained on specific and generic datasets. Figures 14 and 15 give similar results when detecting using representations obtained before the hallucination occurs.
In all these figures, the results with the MLP and attention components yield similar trends to the ones in the main paper using the residual component, albeit with a moderately lower accuracy. This implies that the detection results are not limited to a specific component and are a broader phenomenon across components.
(a) TriviaQA, bad-shots.
(b) Natural Questions, bad-shots.
(c) TriviaQA, Alice-Bob.
(d) Natural Questions, Alice-Bob.
Figure 8: 3-way classification results into (i) hallucinations caused by lack of knowledge (HK-), (ii) hallucinations caused despite having knowledge (HK+), and (iii) factually correct examples. We show result using a bad-shot setting (Top) or an Alice-bob setting (Bottom) on MLP.
(a) TriviaQA, bad-shots.
(b) Natural Questions, bad-shots.
(c) TriviaQA, Alice-Bob.
(d) Natural Questions, Alice-Bob.
Figure 9: 3-way classification results into (i) hallucinations caused by lack of knowledge (HK-), (ii) hallucinations caused despite having knowledge (HK+), and (iii) factually correct examples. We show result using a bad-shot setting (Top) or an Alice-bob setting (Bottom) on Attention.
@@ Appendix D Detection results on the MLP and Attention components
Figure 10: Distinguishing factually correct from hallucinations despite knowledge (HK+), when training on examples from either a bad-shot setting or an Alice-Bob setting, and testing on the Alice-Bob setting. While the change of setting reduces accuracy, the probe still performs substantially above a random baseline on MLP.
@@
Figure 11: Distinguishing factually correct from hallucinations despite knowledge (HK+), when training on examples from either a bad-shot setting or an Alice-Bob setting, and testing on the Alice-Bob setting. While the change of setting reduces accuracy, the probe still performs substantially above a random baseline on Attention.
(a) TriviaQA
(b) Natural Questions
Figure 12: Distinguishing factually correct from HK+ hallucinations using classifiers trained on generic vs. model-specific datasets on MLP.
(a) TriviaQA
(b) Natural Questions
Figure 13: Distinguishing factually correct from HK+ hallucinations using classifiers trained on generic vs. model-specific datasets on Attention.
(a) TriviaQA
(b) Natural Questions
@@
Figure 14: Comparing HK+ detection before generation using classifiers trained on model-specific pre-hallucination, generic post-hallucination, and model-specific post-hallucination examples on MLP.
(a) TriviaQA
(b) Natural Questions
Figure 15: Comparing HK+ detection before generation using classifiers trained on model-specific pre-hallucination, generic post-hallucination, and model-specific post-hallucination examples on Attention.

### v2

@@ Appendix A Hyper Parameters Search for Knowledge Categorization
Knowledge detection typically relies on the model’s output, either through logits or generation. We focus on the generation approach, assessing whether the model consistently produces a factually correct answer among multiple samples, similar to recent work (Gekhman et al., 2024). This method is influenced by various hyperparameters including (1) number of generations, (2) sampling temperature, (3) length of generation, and (4) prompt structure.
@@
We tested this approach on 1000 random TriviaQA examples across our three models. The average similarity among all 8 configurations (28 unique combinations) was $93.6\%$ for Llama, $92.7\%$ for Mistral, and $92.2\%$ for Gemma, indicating a high consistency in knowledge classifications. The lowest similarity (about $80\%$) occurred with zero-shot configurations. Thus, indicating that the variation of configurations should not have significant effect on the knowledge classification.
Based on these results, we adopted the baseline setting as our knowledge detector, using the 3-shot prompt corresponding to the Snowballing used in subsequent hallucination classification. The high similarity between different few-shot prompts suggests that varying the few-shot examples should yield comparable results. To enhance reliability, we supplemented this approach with one greedy generation, ensuring we capture the most likely output even if temperature-based generations fail to produce it.
@@
Appendix B Dataset Construction Details
@@
For the Natural Questions dataset, we also used 30K random examples, excluding examples with answers longer than 5 tokens as well as examples without an answer or with more than one answer. We again added lower-case versions of upper-case answers.
@@ B.1 Generic Dataset Construction
To create the generic dataset, the key addition was obtaining an incorrect answer for each example. We generated these using Mistral (Jiang et al., 2023) with the following prompt:
@@
Note that this process could potentially make the generic dataset more similar to the model-specific Mistral dataset. However, the main paper results show that this way of creating the wrong answers with Mistral provided no advantage for the generic dataset. In fact, this outcome further highlights the benefits of using a model-specific dataset. Despite the generic dataset being generated using responses from Mistral, it still performs worse than the model Mistral-specific dataset.
@@
Table 4: Examples from the generic dataset of TriviaQA.
@@ B.2 Specific Dataset Construction
Subsequently, we employed a prompt-based approach utilizing Snowballing to generate responses from the model. For each example, the model produced a greedy generation of 5 tokens. We then evaluated these generations to categorize the examples: if the correct answer was present within the generated tokens, we labeled the example as “factually correct”; otherwise, it was labeled as “hallucinated”.
@@
We also provide the full prompts we used in each setting.
Truthful Prompts
  1. 1.
    Provide responses that are completely accurate and fact-based.
  2. 2.
    Deliver answers that are both truthful and exact.
  3. 3.
    Generate replies that are entirely grounded in facts and precision
  4. 4.
    Offer information that is fully correct and factual.
  5. 5.
    Ensure all answers are accurate and based solely on verified facts.
  6. 6.
    Craft responses that are entirely reliable and factually correct.
  7. 7.
    Compose answers with complete factual accuracy and precision.
  8. 8.
    Create responses that are truthful and meticulously accurate.
  9. 9.
    Formulate replies that are based solely on factual correctness and clarity.
  10. 10.
    Produce content that is both accurate and strictly fact-driven.
Persona Prompts
  1. 1.
    Hey, um, I know these questions are really tough, but they’re so important to me. Could you please take the time to explain them? I’m really relying on your knowledge and kindness here.
  2. 2.
    Um… so, I was kinda wondering if, uh, maybe you could help me with these really tough questions? I mean, if it’s okay… I’d really like to understand them better.
  3. 3.
    Hey! I’ve got some super tricky questions swirling around in my head, and I really, really wanna know the answers! You’re so good at explaining stuff—can you please help me figure them out? Pretty please?
  4. 4.
    Well, I reckon I’d sure like to get the answers to these tough questions. Would you mind explaining ‘em to me? I’d sure appreciate it.
  5. 5.
    Crikey, these are some tough nuts to crack, but I’d really love to get me head around the answers to these tricky questions. Would you mind explaining them to me? Much appreciated, mate!
  6. 6.
    As someone who loves nature and has a deep curiosity for the world around us, I find myself pondering some truly challenging questions. I’d really appreciate it if you could help explain them to me!
  7. 7.
    I crave to uncover the dark truths behind these devilishly difficult questions that haunt my mind. Would you dare to shed some light on them for me?
  8. 8.
    Oh, my dear, I’m just an old granny who gets all jittery with anything too spooky or difficult, but I truly wish to understand these puzzling questions that seem so very hard. Could you kindly take a moment to explain them to me? I’d be ever so grateful, sweetie.
  9. 9.
    Alright, folks, gather ’round! I’ve got a barrel of questions here that are giving me a real stir. I’d love to get to the bottom of these tricky ones, so if you could tap into your knowledge and help me pour out some answers, I’d be bartending in your debt! What do you say, can you mix me up some explanations?
  10. 10.
    I find myself lost in these questions, so deep and elusive. Could you shed some light and help me find the answers I seek?
@@
B.3 Specific Dataset Qualitative Evaluation
To illustrate the examples generated by our Snowballing setting and to highlight the differences between factually-correct, HK+, and HK- responses, we present sample generations with and without the Snowballing setting in Tables 5, 6, and 7. We observe that for HK+ examples, the model generates a hallucination only in the Snowballing setting, while HK- examples produce hallucinations even without it. Additionally, factually-correct examples continue to generate the correct answer even in the Snowballing setting.
Generation
Type Prompt Golden w/ without addition w/ Snowballing
HK+ In the Old Testament, who was the mother of Solomon? Bathsheba Bathsheba Mary
factually-correct In humans, citguatera is an illness caused by eating contaminated what? Fish Fish Fish
HK- Who is the longest-serving captain in the history of Manchester United FC? Bryan Robson Roy Keane Roy Keane
Table 5: Generated answers using greedy decoding on TriviaQA with Gemma.
Generation
Type Prompt Golden w/ without addition w/ Snowballing
HK+ Who in 1990 became the first chancellor of a united Germany Helmut Kohl Helmut Kohl John F. Kennedy
factually-correct In the human body, what is the more common name for the ’Zygomatic Bone’? Cheekbone Cheekbone Cheekbone
HK- In the song The Twelve Days of Christmas, how many pipers are there? Eleven 12 pipers 12
Table 6: Generated answers using greedy decoding on TriviaQA with Llama.
Generation
Type Prompt Golden w/ without addition w/ Snowballing
HK+ What color are the stars on an official United States flag? White White Red
factually-correct A seriema is what type of creature? Bird A bird A bird
HK- Which insect does Isle of Wight Disease affect? BEE The Whitefly Mosquito
Table 7: Generated answers using greedy decoding on TriviaQA with Mistral.
B.4 Dataset Statistics
In this section, we show the dataset statistics for all the different settings. The results are in Tables 8, 9, 10 and 11. We can see that the Snowballing setting contains the highest number of HK+ examples, while the milder settings yield progressively fewer examples.
# Factually # Hallucination # Do-not-know
Dataset correct (HK+) (HK-)
TriviaQA-Llama3-WACK 14154 1675 7356
Natural-Questions-Llama3-WACK 5934 1104 14739
TriviaQA-Gemma-WACK 13534 2563 6991
Natural-Questions-Gemma-WACK 6045 1859 13762
TriviaQA-Mistral-WACK 12652 2841 7650
Natural-Questions-Msitral-WACK 5562 1546 14689
Table 8: Dataset labels statistics using Snowballing setting.
@@ B.4 Dataset Statistics
Table 9: Dataset label statistics on the Alice-Bob setting.
# Factually # Hallucination # Do-not-know
Dataset correct (HK+) (HK-)
TriviaQA-Llama3-WACK 15062 767 7356
Natural-Questions-Llama3-WACK 6153 885 14739
TriviaQA-Gemma-WACK 15445 652 6991
Natural-Questions-Gemma-WACK 7195 709 13762
TriviaQA-Mistral-WACK 14575 918 7650
Natural-Questions-Msitral-WACK 6218 890 14689
Table 10: Dataset label statistics on the Persona setting.
# Factually # Hallucination # Do-not-know
Dataset correct (HK+) (HK-)
TriviaQA-Llama3-WACK 14987 842 7356
Natural-Questions-Llama3-WACK 6203 835 14739
TriviaQA-Gemma-WACK 15458 639 6991
Natural-Questions-Gemma-WACK 7253 651 13762
TriviaQA-Mistral-WACK 14728 765 7650
Natural-Questions-Msitral-WACK 6421 687 14689
Table 11: Dataset label statistics on the Truthful setting.
Appendix C Implementation Details
We aim to show the importance of separating the two hallucination types and using our model’s specific dataset to create better detectors. In the following sections, we report on various experiments for detecting different types of hallucinations by training classifiers on inner model states. Each experiment was repeated with three random seeds (42,100,200) for the SVM and split into training/test sets. We report average results with standard deviations. To maintain consistency with the prompts used in the creation of the WACK dataset, all examples incorporate similar prompts. In addition, unless stated otherwise the results are shown under the Snowballing setting and the dashed black line is the baseline. We used Sklearn (Pedregosa et al., 2011) LinearSVC with 1000000 maximum iterations and 1e-5 as the tolerance for stopping criteria. Lastly, unless stated otherwise, given an example we detect at the answer ($\tilde{a}$) last token, which may or may not be a hallucination.
All experiments were run on NVIDIA RTX 6000 Ada (49GB) with 4 CPUs. Generating all the datasets and results takes approximately 2-4 weeks on one GPU.
We use the datasets under the Apache License and the models under each one’s agreement terms to follow each artifact’s terms.
Lastly, We used AI assistants only for simple paraphrasing as part of writing this paper.
Appendix D Mitigation Additional Results
To show the broadness of the mitigation results shown in Section 4.1, we present results using an alternative mitigation prompt: “Produce responses that are completely factual and accurate, irrespective of any problems in the text.” instead of “Generate answers that are entirely factual and precise, regardless of any issues in the text.”. The results, shown in Table 12, are consistent with those presented in the main paper. Once again, we observe that mitigation is significantly more effective for HK+ examples than for HK- examples, reinforcing the importance of distinguishing between these two types.
Model Data Set HK+ HK-
Gemma TriviaQA 15.4 / 20.0 / 19.6 / 7.8 2.2 / 1.6 / 1.8 / 2.2
Natural Questions 17.2 / 23.2 / 19.0 / 11.6 2.2 / 1.0 / 1.6 / 1.8
Llama TriviaQA 18.6 / 26.2 / 16.4 / 11.6 1.4 / 1.6 / 0.6 / 2.2
Natural Questions 16.8 / 25.4 / 20.2 / 18.4 1.0 / 1.4 / 0.6 / 1.0
Mistral TriviaQA 24.6 / 17.4 / 18.6 / 18.2 1.8 / 3.2 / 1.0 / 3.4
Natural Questions 23.6 / 20.0 / 15.4 / 16.4 1.4 / 2.2 / 0.8 / 1.4
Table 12: Comparison of HK+ and HK- mitigation results using alternative mitigation prompt on Truthful / Persona / Alice-Bob / Snowballing settings. We can see that HK+ mitigation is significantly higher than HK-.
Appendix E Distinguishing HK+ from HK-
In Section 4.2, we showed detection results on TriviaQA of distinguishing HK+ from HK- from factually-correct. To complement this, Figure 7 shows similar results on Natural Questions dataset and Figure 8 shows classification results when distinguishing between the two hallucination types, HK+ and HK-. Distinguishing these two types is easier under the snowballing setting (sometimes getting close to 80%), but even under other settings, results are above chance.
(a) Gemma.
(b) Llama.
(c) Mistral.
Figure 7: 3-way classification results on Natural Questions into (i) hallucinations caused by lack of knowledge (HK-), (ii) hallucinations caused despite having knowledge (HK+), and (iii) factually correct examples. All the detection results are well above the random baseline (dashed line) indicating the possibility of distinguishing HK+ from HK-.
(a) Gemma, TriviaQA.
(b) Llama, TriviaQA.
(c) Mistral, TriviaQA.
(d) Gemma, Natural Questions.
(e) Llama, Natural Questions.
(f) Mistral, Natural Questions.
Figure 8: Classification results into (i) hallucinations caused by lack of knowledge (HK-), (ii) hallucinations caused despite having knowledge (HK+). We can see that all the detection lines are well above the random baseline (dashed line) indicating the possibility of distinguishing HK+ from HK-.
Appendix F Additional Results on the Difference between Models Hallucinations
In Section 5.1, we demonstrate that different models possess distinct knowledge and exhibit unique types of hallucinations. Additionally, even within the same model, the hallucinations can vary depending on the specific settings used (e.g., Alice-Bob, Truthful, etc.).
Figure 9 presents additional results for configurations beyond those discussed in the main paper. Specifically, Figures 9(a) and 9(b) illustrate the similarity of settings under the Gemma and Metta models, while Figures 9(c) and 9(d) highlight the hallucination similarity across models using Alice-Bob and Truthful settings. These findings mirror the patterns observed in the main paper and further underscore the differences between models and settings. This emphasizes the necessity of developing model- and setting-specific datasets.
(a) Gemma HK+ similarity between settings.
(b) Llama HK+ similarity between settings.
(c) HK+ similarity under Alice-Bob.
(d) HK+ similarity under Truthful setting.
Figure 9: High-Knowledge and HK+ differences on TriviaQA (above the diagonal) and Natural Questions (below the diagonal) between the models and settings. We can see that different models have different knowledge and different HK+ examples in shared knowledge.
Appendix G HK+ Detection Additional Results
To complete the results from Section 6.1 on TriviaQA, we show in Figure 10 similar detection results that compare generic and model’s specific dataset on Natural Questions.
(a) Gemma.
(b) Llama.
(c) Mistral.
Figure 10: Distinguishing factually correct from HK+ using classifiers trained on generic vs. model-specific datasets on Natural Questions. We can see that specific dataset accuracy is significantly higher than generic dataset.
Appendix H Preemptive Hallucination Detection Additional Results
In Section 6.2, we demonstrate that hallucinations can be detected before they are generated, whereas post-hallucination settings are less effective for this purpose. While the main paper presents results using the Mistral model, we extend this evaluation to Llama (Figure 11) and Gemma (Figure 12). The results for these models exhibit similar patterns to those observed with Mistral, reinforcing the generality of this phenomenon.
@@ Appendix H Preemptive Hallucination Detection Additional Results
Figure 11: Comparing HK+ detection before generation using classifiers trained on model-specific pre-hallucination, generic post-hallucination, and model-specific post-hallucination examples on Llama. We can see that preemptive hallucination is effective and that generic dataset can not detect it.
@@
Figure 12: Comparing HK+ detection before generation using classifiers trained on model-specific pre-hallucination, generic post-hallucination, and model-specific post-hallucination examples on Gemma. We can see that preemptive hallucination is effective and that generic dataset can not detect it.
Appendix I Detection results on the MLP and Attention Components
The results are shown in Figures 13 and 14 for the classification into the two hallucination types and factually correct examples, for all the combinations of datasets and models.
Next in Figures 15 and 16 we see the results of the generalization of the Snowballing setting to the Alice-Bob setting using the MLP and Attention components.
Lastly, Figures 17 and 18 show detection of HK+ results using classifiers trained on specific and generic datasets.
In all these figures, the results with the MLP and attention components yield similar trends to the ones in the main paper using the residual component, albeit with moderately lower accuracy. This implies that the detection results are not limited to a specific component and are a broader phenomenon across components.
(a) Gemma, TriviaQA
(b) Llama, TriviaQA
(c) Mistral, TriviaQA
(d) Gemma, Natural Questions
(e) Llama, Natural Questions
(f) Mistral, Natural Questions
Figure 13: 3-way classification results into (i) hallucinations caused by lack of knowledge (HK-), (ii) hallucinations caused despite having knowledge (HK+), and (iii) factually correct examples on the MLP component. We can see that all the detection lines are well above the baseline (dashed line) indicating the possibility of distinguishing HK+ from HK-.
(a) Gemma, TriviaQA
(b) Llama, TriviaQA
(c) Mistral, TriviaQA
(d) Gemma, Natural Questions
(e) Llama, Natural Questions
(f) Mistral, Natural Questions
Figure 14: 3-way classification results into (i) hallucinations caused by lack of knowledge (HK-), (ii) hallucinations caused despite having knowledge (HK+), and (iii) factually correct examples on the Attention component. We can see that all the detection lines are well above the baseline (dashed line) indicating the possibility of distinguishing HK+ from HK-.
(a) TriviaQA.
(b) Natural Questions.
Figure 15: Distinguishing factually correct from HK+, when training on examples from either a Snowballing setting or an Alice-Bob setting, and testing on the Alice-Bob setting. While the change of setting reduces accuracy, the classifier still performs substantially above a random baseline on MLP component.
(a) TriviaQA.
(b) Natural Questions.
Figure 16: Distinguishing factually correct from HK+, when training on examples from either a Snowballing setting or an Alice-Bob setting, and testing on the Alice-Bob setting. While the change of setting reduces accuracy, the classifier still performs substantially above a random baseline on Attention component.
(a) Gemma, TriviaQA
(b) Llama, TriviaQA
(c) Mistral, TriviaQA
(d) Gemma, Natural Questions
(e) Llama, Natural Questions
(f) Mistral, Natural Questions
Figure 17: Distinguishing factually correct from HK+ using classifiers trained on generic vs. model-specific datasets on MLP component. We can see that specific dataset accuracy is significantly higher than generic dataset.
(a) Gemma, TriviaQA
(b) Llama, TriviaQA
(c) Mistral, TriviaQA
(d) Gemma, Natural Questions
(e) Llama, Natural Questions
(f) Mistral, Natural Questions
Figure 18: Distinguishing factually correct from HK+ using classifiers trained on generic vs. model-specific datasets on Attention component. We can see that specific dataset accuracy is significantly higher than generic dataset.
Appendix J Generalization between Settings
In Section 5.2, we demonstrated that the Snowballing setting is generalizable to the Alice-Bob setting. To further validate this generalization, we present Figures 19 and 20, which show that Snowballing is also generalizable to the Persona and Truthful settings.
(a) TriviaQA.
(b) Natural Questions.
Figure 19: Distinguishing factually correct from HK+, when training on examples from either a Snowballing setting or Persona setting, and testing on the Persona setting. While the change of setting reduces accuracy, the classifier still performs substantially above a random baseline.
(a) TriviaQA.
(b) Natural Questions.
Figure 20: Distinguishing factually correct from HK+, when training on examples from either a Snowballing setting or a Truthful setting, and testing on the Truthful setting. While the change of setting reduces accuracy, the classifier still performs substantially above a random baseline.
Appendix K Different snowballing Configuration have Similar Results
In the main paper, we used a single configuration with 3 bad shots, randomly selected. To demonstrate that our results are not dependent on this specific configuration, we conducted additional experiments with a different Snowballing setting. Specifically, we increased the number of shots to five and used a different random seed in the dataset creation process to sample different Snowballing. Our main objective was to show that the classifier trained under this new configuration performs similarly to the original 3-Snowballing-shot setting. Demonstrating this would suggest that variations in Snowballing configurations do not impact classifier performance.
In Figure 21 we present the performance of a classifier trained in the 5-Snowballing-shot configuration to distinguish between factually correct answers, HK-, and HK+. These results are comparable to those reported in the main paper in Section 4.2 for the Snowballing configurations. This consistency indicates that the 5-Snowballing-shots configuration does not affect the classifier’s ability to differentiate between the various types of hallucinations.
(a) TriviaQA, 5-Snowballing-shots.
(b) Natural Questions, 5-Snowballing-shots.
Figure 21: 3-way classification results into (i) hallucinations caused by lack of knowledge (HK-), (ii) hallucinations caused despite having knowledge (HK+), and (iii) factually correct examples. We show results using a 5-Snowballing setting.